# CTPU-PTC-23-29

### v1

@@ Abstract
We show that the axion-$U(1)$ model, for a finite amount of rolling, potentially explains the deviation from astrophysical background in PTA data, and is consistent with the given spectral shape, together with interesting phenomenological signatures such as chiral primordial GW, scalar induced GW and PBH production. We find that such a parameter region is not excluded but is close to the perturbativity and backreaction bounds derived in small/mild backreaction regime, and thereby there is a need for non-linear analysis. We also note that there is a clear smoking gun of axion inflation coupled to gauge fields, namely statistical parity violation, hence it is expected that the resultant GW background is almost perfectly chiral 555possibly with a scale dependent chirality [92], which is a concrete prediction for forthcoming surveys [93, 94, 95, 96, 97].
Acknowledgement. We are grateful to Robert Caldwell, Angelo Caravano, Emanuela Dimastrogiovanni, Matteo Fasiello, and Daniel G. Figueroa for the comments on the draft. C.U. dedicates this work to Harun Kolçak, and thanks his family for their support. C.U. is supported by the Kreitman fellowship of the Ben-Gurion University, and the Excellence fellowship of the Israel Academy of Sciences and Humanities, and the Council for Higher Education. A.P. is supported by IBS under the project code, IBS-R018-D1. I.O. is supported by JSPS KAKENHI Grant No. JP20H05859 and 19K14702.

### v3

@@ Abstract
We show that the axion-$U(1)$ model, for a finite amount of rolling, potentially explains the deviation from astrophysical background in PTA data, and is consistent with the given spectral shape, together with interesting phenomenological signatures such as chiral primordial GW, scalar induced GW and PBH production. We find that such a parameter region is not excluded but is close to the perturbativity and backreaction bounds derived in small/mild backreaction regime, and thereby there is a need for non-linear analysis. We also note that there is a clear smoking gun of axion inflation coupled to gauge fields, namely statistical parity violation, hence it is expected that the resultant GW background is almost perfectly chiral 888possibly with a scale dependent chirality [120], which is a concrete prediction for forthcoming surveys [121, 122, 123, 124, 125].
Acknowledgement. We are grateful to Robert Caldwell, Angelo Caravano, Emanuela Dimastrogiovanni, Matteo Fasiello, Daniel G. Figueroa and Eiichiro Komatsu for the comments on the draft. C.U. thanks Yann Gouttenoire and Lorenzo Sorbo for the discussions on effective degrees of freedom from gauge fields. C.U. dedicates this work to Harun Kolçak, and thanks his family for their support. C.U. is supported by the Kreitman fellowship of the Ben-Gurion University, and the Excellence fellowship of the Israel Academy of Sciences and Humanities, and the Council for Higher Education. A.P. is supported by IBS under the project code, IBS-R018-D1. I.O. is supported by JSPS KAKENHI Grant No. JP20H05859 and 19K14702.